# A Large-scale Empirical Study on Improving the Fairness of Deep Learning Models

### 08 Jan 2024 v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; January 2024; price: 15.00isbn: 978-1-4503-XXXX-X/18/06copyright: none
@@ 1. Introduction
In recent years, Artificial Intelligence (AI) systems and Deep Learning (DL) models have attracted significant attention for their remarkable performance across diverse application domains, such as image processing (Saarinen, 1994; González and Woods, 2008), machine translation (Brown et al., 1993; Koehn, 2005; Chiang, 2005; Koehn et al., 2007) and software engineering (Jr., 1987; Royce, 1987). However, the increasing adoption of DL models in real practice also poses new challenges to them. In particular, they have displayed troubling discriminatory tendencies in certain applications, such as elevated error rates, when confronted with specific groups or populations defined by attributes deemed protected or sensitive (Friedler et al., 2019; Mehrabi et al., 2022; Zhang and Sun, 2022; Biswas and Rajan, 2020; Chen et al., 2022b; Chakraborty et al., 2019), including factors like race, gender, and age, etc., causing ethical or even safety issues that will largely affect the reliability and usability of them in real practice. For example, AI judge (Steinitz, 2014; Green and Chen, 2019; Huq, 2018; Xu, 2022) may exhibit biased decision-making, leading to unequal treatment of individuals based on their sensitive attributes, such as race. This bias can result in disparities in sentencing or legal outcomes, raising concerns about fairness and justice within the legal system. Similarly, AI-powered hiring (Palshikar et al., 2018; Harirchian et al., 2022; Roemmich et al., 2023) and loan systems (van Zetten et al., 2022; Galindo and Tamayo, 2000; Wang et al., 2011) tend to favor applicants of a particular gender or race. In other words, the manifestation of unfair behaviors in AI systems and DL models has raised profound ethical concerns, transforming fairness into an essential ethical and often legally mandated prerequisite for the widespread adoption and utilization of DL models across various real-world contexts.
@@
To improve the fairness of DL models, many approaches have been proposed in the last decades, particularly in the application domain of image classification, which has been widely adopted in diverse real-world applications. Typically, existing approaches can be classified into three categories according to the stage where de-biasing is applied, i.e., pre-processing, in-processing, and post-processing. Pre-processing techniques (Buda et al., 2018; Japkowicz and Stephen, 2002; Qraitem et al., 2022; Yao et al., 2022; Sagawa et al., 2020; Ramaswamy et al., 2021; Wang et al., 2020) endeavor to mitigate model biases by optimizing the training data, such as performing data transformation  (Yao et al., 2022), augmentation  (Ramaswamy et al., 2021), sampling  (Buda et al., 2018; Japkowicz and Stephen, 2002; Sagawa et al., 2020; Wang et al., 2020; Qraitem et al., 2022) and so on. In contrast, in-processing techniques (Tzeng et al., 2015; Hong and Yang, 2021; Sarridis et al., 2023; Jung et al., 2021; Kehrenberg et al., 2020) dedicate to improve the model fairness by optimizing the model training process, such as updating the training loss (Hong and Yang, 2021; Sarridis et al., 2023) or model structures (Jung et al., 2021; Kehrenberg et al., 2020). Different from the former two categories, post-processing techniques (Kim et al., 2019) aim to make the deployed models produce fairer outputs without changing the model structures and parameters, such as optimizing the given inputs (Zhang et al., 2022b; Wang et al., 2022). Since they do not change the original models, they can be flexibly applied to improve the fairness of deployed models in real practice.
@@
    Existing methods tend to be insensitive to different evaluation metrics, and thus subset of the metrics can be representative in future studies.
@@
    Existing methods are sensitive to different experimental datasets, and thus the conclusions over a certain dataset may not be generalizable to others.
@@ 2. Background
Firstly, we would like to clarify the problem and concept of fairness issues in deep learning models. Furthermore, we will also conduct a succinct literature review of the most recent research on fairness improvement and identify the limitations of them, which motivates the necessity and signficance of this study.
@@ 2.1. Fairness Issues
In fact, software fairness issues (i.e., unfair software predictions) (Finkelstein et al., 2008) have been a growing area of concern in both the Software Engineering (SE) and Deep Learning (DL) research communities. For example, they are typically referred to as fairness defects  (Brun and Meliou, 2018) or fairness bugs  (Chen et al., 2022b) in the SE community. In this study, we refer to the unfair predictions produced by DL models as fairness issues by following existing work (Finkelstein et al., 2008), indicating a discordance between existing and expected fairness criteria. Many testing (Chen et al., 2022a; Galhotra et al., 2017; Patel et al., 2022; Guo et al., 2023; Aggarwal et al., 2019; Soremekun et al., 2022) and repair studies (Nguyen et al., 2023; Salimi et al., 2019; Zhang et al., 2022a; Sun et al., 2022) aim to automatically identify and fix such issues. In this study, we direct our attention specifically to the techniques that aim to resolve the fairness issues in image classification models since they have been already widely deployed in practice as explained in the introduction. In addition, this study focuses on the group fairness of DL models.
@@
Image FairHSIC  (Quadrianto et al., 2019), US  (Sagawa et al., 2020), OS  (Wang et al., 2020), GAN-debiasing  (Ramaswamy et al., 2021), CGL  (Jung et al., 2022), UW  (Qraitem et al., 2022), BM  (Qraitem et al., 2022) gender, background, age, race, color/grayscale EOD, DEO, BA, BC, KL
In-processing Tabular META  (Celis et al., 2019), AD  (Zhang et al., 2018), GR  (Agarwal et al., 2018), MAAT  (Chen et al., 2022b), PR  (Kamishima et al., 2012), FMT  (Jiang et al., 2023), CARE  (Sun et al., 2022), Parfait-ML  (Tizpaz-Niari et al., 2022) gender, age, race NMI, PI/MI, NPI, UEI, CVS, SPD, AAOD, EOD
Image Adv  (Wang et al., 2020), DI  (Wang et al., 2020), BC+BB  (Hong and Yang, 2021), MFD  (Jung et al., 2021), FairBatch  (Roh et al., 2021), FSCL  (Park et al., 2022), FairGRAPE  (Lin et al., 2022), FDR  (Mao et al., 2023), FLAC  (Sarridis et al., 2023) background color, color (bright/dark), age, race, texture bias, gender, color/grayscale EOD, BC, WA, DEO, SPD, AED
@@
Image FairReprogram  (Zhang et al., 2022b), Multiaccuracy  (Kim et al., 2019), FAAP  (Wang et al., 2022) gender DEO, SPD
@@ 2.2. Limitations in Existing Studies
In recent years, there has been a growing focus among researchers on the fairness issues, leading to various of fairness improvement methods being proposed. In order to gain a deeper understanding of existing studies, we have conducted an succinct literature review of relevant papers published recently in the prominent conferences and journals in the Software Engineering (SE) and Artificial Intelligence (AI) domains, including ICSE, ESEC/FSE, TOSEM, AIES, ICLR, ECCV, CVPR, NeurIPS, etc. Table 1 presents the details of these methods that aim to improve DL models’ fairness.
In this table, besides the essential information of the method (e.g., method names and categories), we further classify the methods into Tabular and Image based on the type of dataset they are designed for. Additionally, we display the union sets of sensitive attributes and fairness metrics adopted by all methods within each category. After a careful examination, we have the following two major observations regarding the limitations in existing studies.
@@
(Inconsistent Fairness Evaluation Metrics): Most of the existing studies only adopted only one or two metrics, which may not offer a complete assessment of model fairness. Moreover, some works introduce unique fairness metrics (Kim et al., 2019; Lin et al., 2022) tailored to their specific task scenarios, but these metrics are rarely cited or adopted by other research, raising questions about the reliability and generalizability of these metrics. In particular, the different metrics adopted by existing studies further hinder the comparison of different techniques.
The limitations above pose a emergent need to conduct a large-scale empirical study for comprehensively comparing their performance under a uniform context, which motivates this work. In particular, to address these limitations, we have replicated the experimental results of many existing approaches over their original datasets and then further evaluated their effectiveness over three different datasets involving both facial and non-facial images (i.e., objects). Furthermore, we have employed all the widely-used performance metrics (i.e., five fairness metrics and two accuracy metrics) for systematically demonstrating the effectiveness of them. In this way, we can better understand the performance of existing techniques and derive valuable insights for bias mitigation in deep learning models. Ultimately, our aspiration is to provide meaningful guidance that will facilitate more effective explorations of fairness issues in the realm of image fairness.
@@ 3. Studied Methods in This Study
As summarized in Table 1, there are many novel methods have been published in a wide range of conferences and journals, aiming to enhance the fairness of neural networks. The latest research (Chen et al., 2024) has explored the performance of existing methods that focus on tabular datasets, while those methods targeting image datasets are still not systematically investigated. Therefore, our study will primarily focus on fairness improving methods for image datasets. Specifically, we follow the selection criteria below to select the methods ultimately studied in our study.
@@
(Latest Studies): To ensure the effectiveness and representativeness of our study, we prefer to select the methods that were most recently proposed since they usually can achieve much better performance than the previous.
As a result, 13 state-of-the-art methods (including two variants of FR (Zhang et al., 2022b)) were included, involving all the three categories, i.e., pre-processing, in-processing, and post-processing. We also emphasized these methods in bold in Table 1 and introduced them as follows.
@@ 3.1. Pre-processing
Pre-processing methods, as a general practice, are primarily concerned with reducing discrimination and bias within the training data, with the ultimate objective of improving the fairness of the resulting trained model. These techniques involve data transformations and augmentations designed to mitigate inherent forms of discrimination. Pre-processing strategies are employed when the algorithm has the capacity to modify the training data to mitigate bias. From the plethora of available pre-processing methods  (Buda et al., 2018; Japkowicz and Stephen, 2002; Wang et al., 2020; Byrd and Lipton, 2019; Qraitem et al., 2022; Jung et al., 2022; Yao et al., 2022; Quadrianto et al., 2019; Sagawa et al., 2020), we specifically concentrate on the following four approaches as the representative based on the selection criteria.
@@
  • Oversampling (OS)  (Wang et al., 2020) is employed to rectify class imbalance by duplicating samples, thereby equalizing class sizes. This approach has been adapted to effectively address bias by ensuring balance within sensitive subgroups. In practice, we ascertain the size of the largest subgroup and then replicate samples in other subgroups proportionally. However, it’s important to exercise caution, as excessive sample duplication can result in overfitting, especially when dealing with highly parameterized models like deep neural networks.
@@ 3.2. In-processing
  • Domain Independent Training (DI)  (Wang et al., 2020; Qraitem et al., 2022) involves the utilization of extra prediction heads for each bias group. In response to the issue of discriminative models learning unnecessary domain-specific class boundaries, DI promotes the adoption of distinct domain-specific classifiers. Nonetheless, training separate classifiers diminishes data exposure. To address this concern, DI proposes a shared feature representation alongside an ensemble of classifiers. It recommends direct reasoning on class boundaries of domains, effectively eradicating class-domain correlations.
@@
  • MMD-based Fair Distillation (MFD)  (Jung et al., 2021) combines feature distillation with maximum mean discrepancy (MMD) to enhance both prediction accuracy and fairness simultaneously. The method’s theoretical foundation demonstrates that its MMD-based regularization fosters fairness by aligning group-conditioned features of the student model across sensitive groups and aligning them with the teacher model’s group-averaged features, resulting in improved accuracy as well.
@@ 3.3. Post-processing
  • FairReprogram (FR)  (Zhang et al., 2022b) leverages model reprogramming techniques to address scenarios where it’s not feasible to modify existing models. Instead, it appends a set of perturbations known as the ”fairness trigger” to the input data, optimizing it in a min-max framework to align with fairness criteria. In particular, FR applies two methodologies to append the trigger, corresponding to its 2 variants, named FR-P (appending the trigger like a patch to the original image) and FR-B (appending the trigger at the border of the original image). We applied both of them in our study. Specifically, they perturb the inputs by adding a constant global vector/feature, which is tailored to improve fairness while keeping the deep learning model unchanged. FairReprogram is a versatile framework applicable to various tasks and domains, making it a generic solution.
  • Fairness-Aware Adversarial Perturbation (FAAP)  (Wang et al., 2022) mitigates unfairness in deployed deep models by updating the given input without modifying model parameters or structures. Specifically, it achieves this by learning to perturb input data, effectively blinding deployed models on fairness-related attributes such as gender or ethnicity. FAAP employs a discriminator-generator adversarial framework to ensure that protected attributes are not correlated with model predictions, making it a flexible and practical approach to addressing fairness bugs in real-world AI systems.

### 24 Mar 2024 v2

# A Large-Scale Empirical Study on Improving the Fairness of Image Classification Models

@@ Abstract.
Empirical study, Deep learning, Model fairness
copyright: acmlicenseddoi: 10.1145/3650212.3652122journalyear: 2024submissionid: issta24main-p445-pisbn: 979-8-4007-0612-7/24/09conference: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis; September 16–20, 2024; Vienna, Austriabooktitle: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA ’24), September 16–20, 2024, Vienna, Austriaccs: Computing methodologies Computer visionccs: Software and its engineering Software reliabilityccs: General and reference Empirical studies
@@ 1. Introduction
In recent years, Artificial Intelligence (AI) systems and Deep Learning (DL) models have attracted significant attention for their remarkable performance across diverse application domains, such as image processing (Saarinen, 1994; González and Woods, 2008), machine translation (Brown et al., 1993; Koehn, 2005; Chiang, 2005; Koehn et al., 2007) and software engineering (Jr., 1987; Royce, 1987; Yang et al., 2023; You et al., 2023). However, the increasing adoption of DL models in real practice also poses new challenges to them. In particular, they have displayed troubling discriminatory tendencies in certain applications, such as elevated error rates, when confronted with specific groups or populations defined by attributes deemed protected or sensitive (Friedler et al., 2019; Mehrabi et al., 2022; Zhang and Sun, 2022; Biswas and Rajan, 2020; Chen et al., 2022b; Chakraborty et al., 2019), including factors like race, gender, and age, etc., causing ethical or even safety issues that will largely affect the reliability and usability of them in real practice. For example, AI judge (Steinitz, 2014; Green and Chen, 2019; Huq, 2018; Xu, 2022) may exhibit biased decision-making, leading to unequal treatment of individuals based on their sensitive attributes, such as race. This bias can result in disparities in sentencing or legal outcomes, raising concerns about fairness and justice within the legal system. Similarly, AI-powered hiring (Palshikar et al., 2018; Harirchian et al., 2022) and loan systems (Galindo and Tamayo, 2000; Wang et al., 2011) tend to favor applicants of a particular gender or race. In other words, the manifestation of unfair behaviors in AI systems and DL models has raised profound ethical concerns, transforming fairness into an essential ethical and often legally mandated prerequisite for the widespread adoption and utilization of DL models across various real-world contexts.
@@
To improve the fairness of DL models, many approaches have been proposed in the last decades, particularly in the application domain of image classification, which has been widely adopted in diverse real-world applications. Typically, existing approaches can be classified into three categories according to the stage where de-biasing is applied, i.e., pre-processing, in-processing, and post-processing. Pre-processing techniques (Buda et al., 2018; Japkowicz and Stephen, 2002; Qraitem et al., 2022; Yao et al., 2022; Sagawa et al., 2020; Ramaswamy et al., 2021; Wang et al., 2020a) endeavor to mitigate model biases by optimizing the training data, such as performing data transformation  (Yao et al., 2022), augmentation  (Ramaswamy et al., 2021), sampling  (Buda et al., 2018; Japkowicz and Stephen, 2002; Sagawa et al., 2020; Wang et al., 2020a; Qraitem et al., 2022) and so on. In contrast, in-processing techniques (Tzeng et al., 2015; Hong and Yang, 2021; Sarridis et al., 2023; Jung et al., 2021; Kehrenberg et al., 2020) dedicate to improve the model fairness by optimizing the model training process, such as updating the training loss (Hong and Yang, 2021; Sarridis et al., 2023) or model structures (Jung et al., 2021; Kehrenberg et al., 2020). Different from the former two categories, post-processing techniques (Kim et al., 2019) aim to make the deployed models produce fairer outputs without changing the model structures and parameters, such as optimizing the given inputs (Zhang et al., 2022; Wang et al., 2022). Since they do not change the original models, they can be flexibly applied to improve the fairness of deployed models in real practice.
@@
    Existing methods tend to be insensitive to different evaluation metrics, and thus a subset of the metrics can be representative in future studies.
@@
    Existing methods are sensitive to different experimental datasets, and thus the conclusions drawn from a certain dataset may not be generalizable to others.
@@ 2. Background
Firstly, we would like to clarify the problem and concept of fairness issues in deep learning models. Furthermore, we will also conduct a succinct literature review of the most recent research on fairness improvement and identify the limitations of them, which motivates the necessity and significance of this study.
@@ 2.1. Fairness Issues
In fact, software fairness issues (i.e., unfair software predictions) (Finkelstein et al., 2008) have been a growing area of concern in both the Software Engineering (SE) and Deep Learning (DL) research communities. For example, they are typically referred to as fairness defects  (Brun and Meliou, 2018) or fairness bugs  (Chen et al., 2022b) in the SE community. In this study, we refer to the unfair predictions produced by DL models as fairness issues by following existing work (Finkelstein et al., 2008), indicating a discordance between existing and expected fairness criteria. Many testing (Chen et al., 2022a; Galhotra et al., 2017; Patel et al., 2022; Guo et al., 2023; Aggarwal et al., 2019; Sun et al., 2020; Soremekun et al., 2022; Sun et al., 2022b; Wang et al., 2020b; Shen et al., 2022) and repair studies (Nguyen et al., 2023; Salimi et al., 2019; Sun et al., 2022a) aim to automatically identify and fix such issues. In this study, we direct our attention specifically to the techniques that aim to resolve the fairness issues in image classification models since they have been already widely deployed in practice as explained in the introduction. In addition, this study focuses on the group fairness of DL models.
@@
Image FairHSIC  (Quadrianto et al., 2019), US  (Sagawa et al., 2020), OS  (Wang et al., 2020a), UW  (Qraitem et al., 2022), GAN-debiasing  (Ramaswamy et al., 2021), CGL (Jung et al., 2022), BM  (Qraitem et al., 2022) gender, background, age, race, color/grayscale EOD, DEO, BA, BC, KL
In-processing Tabular META  (Celis et al., 2019), AD  (Zhang et al., 2018), GR  (Agarwal et al., 2018), MAAT  (Chen et al., 2022b), PR  (Kamishima et al., 2012), FMT  (Jiang et al., 2023), CARE  (Sun et al., 2022a), Parfait-ML  (Tizpaz-Niari et al., 2022) gender, age, race NMI, PI/MI, NPI, UEI, CVS, SPD, AAOD, EOD
Image Adv  (Wang et al., 2020a), DI  (Wang et al., 2020a), BC+BB  (Hong and Yang, 2021), MFD  (Jung et al., 2021), FairBatch  (Roh et al., 2021), FSCL  (Park et al., 2022), FairGRAPE  (Lin et al., 2022), FDR  (Mao et al., 2023), FLAC  (Sarridis et al., 2023) background color, color (bright/dark), age, race, texture bias, gender, color/grayscale EOD, BC, WA, DEO, SPD, AED
@@
Image FairReprogram  (Zhang et al., 2022), Multiaccuracy  (Kim et al., 2019), FAAP  (Wang et al., 2022) gender DEO, SPD
@@ 2.2. Limitations in Existing Studies
In recent years, there has been a growing focus among researchers on fairness issues, leading to various of fairness improvement methods being proposed. In order to gain a deeper understanding of existing studies, we have conducted a succinct literature review of relevant papers published recently in prominent conferences and journals in the Software Engineering (SE) and Artificial Intelligence (AI) domains, including ICSE, ESEC/FSE, TOSEM, AIES, ICLR, ECCV, CVPR, NeurIPS, etc. Table 1 presents the details of these methods that aim to improve DL models’ fairness.
In this table, besides the essential information of the method (e.g., method names and categories), we further classify the methods into Tabular and Image based on the type of dataset they are designed for. Additionally, we display the union sets of sensitive attributes and fairness metrics adopted by all methods within each category. Please note that there are also different classification methods (Caton and Haas, 2020). In this paper we adopt the most widely-used one  (Mehrabi et al., 2022; Zhang and Sun, 2022; Biswas and Rajan, 2020) for easy understanding. After a careful examination, we have the following two major observations regarding the limitations in existing studies.
@@
(Inconsistent Fairness Evaluation Metrics): Most of the existing studies only adopted one or two metrics, which may not offer a complete assessment of model fairness. Moreover, some works introduce unique fairness metrics (Kim et al., 2019; Lin et al., 2022) tailored to their specific task scenarios, but these metrics are rarely cited or adopted by other research, raising questions about the reliability and generalizability of these metrics. In particular, the different metrics adopted by existing studies further hinder the comparison of different techniques.
The limitations above pose an emergent need to conduct a large-scale empirical study for comprehensively comparing their performance under a uniform context, which motivates this work. In particular, to address these limitations, we have replicated the experimental results of many existing approaches over their original datasets and then further evaluated their effectiveness over three different datasets involving both facial and non-facial images (i.e., objects). Furthermore, we have employed all the widely-used performance metrics (i.e., five fairness metrics and two accuracy metrics) for systematically demonstrating the effectiveness of them. In this way, we can better understand the performance of existing techniques and derive valuable insights for bias mitigation in deep learning models. Ultimately, our aspiration is to provide meaningful guidance that will facilitate more effective explorations of fairness issues in the realm of image fairness.
@@ 3. Studied Methods in This Study
As summarized in Table 1, many novel methods have been published in a wide range of conferences and journals, aiming to enhance the fairness of neural networks. The latest research (Chen et al., 2024) has explored the performance of existing methods that focus on tabular datasets, while those methods targeting image datasets have not been systematically investigated yet. Therefore, our study will primarily focus on fairness improving methods for image datasets. Specifically, we follow the selection criteria below to select the methods ultimately studied in our study.
@@
(Latest Studies): To ensure the effectiveness and representativeness of our study, we prefer to select the methods that were most recently proposed since they usually can achieve much better performance than the previous ones.
As a result, 13 state-of-the-art methods (including two variants of FR (Zhang et al., 2022)) were included, involving all the three categories, i.e., pre-processing, in-processing, and post-processing. We also emphasized these methods in bold in Table 1 and introduced them as follows.
@@ 3.1. Pre-processing
Pre-processing methods, as a general practice, are primarily concerned with reducing discrimination and bias within the training data, with the ultimate objective of improving the fairness of the resulting trained model. These techniques involve data transformations and augmentations designed to mitigate inherent forms of discrimination. Pre-processing strategies are employed when the algorithm has the capacity to modify the training data to mitigate bias. From the plethora of available pre-processing methods  (Buda et al., 2018; Japkowicz and Stephen, 2002; Wang et al., 2020a; Byrd and Lipton, 2019; Qraitem et al., 2022; Jung et al., 2022; Yao et al., 2022; Quadrianto et al., 2019; Sagawa et al., 2020), we specifically concentrate on the following four approaches as the representative based on the selection criteria.
@@
  • Oversampling (OS)  (Wang et al., 2020a) is employed to rectify class imbalance by duplicating samples, thereby equalizing class sizes. This approach has been adapted to effectively address bias by ensuring balance within sensitive subgroups. In practice, we ascertain the size of the largest subgroup and then replicate samples in other subgroups proportionally. However, it’s important to exercise caution, as excessive sample duplication can result in overfitting, especially when dealing with highly parameterized models like deep neural networks.
@@ 3.2. In-processing
  • Domain Independent Training (DI)  (Wang et al., 2020a; Qraitem et al., 2022) involves the utilization of extra prediction heads for each bias group. In response to the issue of discriminative models learning unnecessary domain-specific class boundaries, DI promotes the adoption of distinct domain-specific classifiers. Nonetheless, training separate classifiers diminishes data exposure. To address this concern, DI proposes a shared feature representation alongside an ensemble of classifiers. It recommends direct reasoning on class boundaries of domains, effectively eradicating class-domain correlations.
@@
  • MMD-based Fair Distillation (MFD)  (Jung et al., 2021) combines feature distillation with maximum mean discrepancy (MMD) to enhance both prediction accuracy and fairness simultaneously. The theoretical foundation demonstrates that its MMD-based regularization fosters fairness by aligning group-conditioned features of the student model across sensitive groups and aligning them with the teacher model’s group-averaged features, resulting in improved accuracy as well.
@@ 3.3. Post-processing
  • FairReprogram (FR)  (Zhang et al., 2022) leverages model reprogramming techniques to address scenarios where it’s not feasible to modify existing models. Instead, it appends a set of perturbations known as the ”fairness trigger” to the input data, optimizing it in a min-max framework to align with fairness criteria. In particular, FR applies two methodologies to append the trigger, corresponding to its 2 variants, named FR-P (appending the trigger like a patch to the original image) and FR-B (appending the trigger at the border of the original image). We applied both of them in our study. Specifically, they perturb the inputs by adding a constant global vector/feature, which is tailored to improve fairness while keeping the deep learning model unchanged. FairReprogram is a versatile framework applicable to various tasks and domains, making it a generic solution.
  • Fairness-Aware Adversarial Perturbation (FAAP)  (Wang et al., 2022) mitigates unfairness in deployed deep models by updating the given input without modifying model parameters or structures. Specifically, it achieves this by learning to perturb input data, effectively blinding deployed models on fairness-related attributes such as gender or ethnicity. FAAP employs a discriminator-generator adversarial framework to ensure that protected attributes are not correlated with model predictions, making it flexible and practical to addressing fairness bugs in real-world AI systems.